OpenStudy (anonymous):

a sample measures 2.00 L of stp. calculate in Cif the volume changes to 1.00 L and the pressure is changed to .500 atm

OpenStudy (anonymous):

do you mean in CC (Cubic Centimeter) and i s the temperature constant assuming temperature to be constant.... \[P _{1} V _{1}=P _{2} V _{2}\] THEREFORE...\[X*2=.500*1\] X=0.25 L (1 LITER=1000CC) FINAL VOLUME IS 250CC

OpenStudy (jfraser):

I think she means C as in new temp, since BOTH the volume and pressure changes are given. In this case, you should use the combined gas law:\[\frac{P_1V_1}{T_1} = \frac{P_2V_2}{T_2}\]and solve for the new temp.